3. Solve. f(x) = 2x* + x3 – 19x? -9x+9
Answer #1

21- 94 - 198 - و ولا و132 18-12-13 - 3عدد - 22 = 2x1.. =1141) - 1741) - 187(+1) + 2x+1) (و + 184 -1-2) (at) = = (+1[ - (3-1)